Circular No. 8085

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003bj IN IC 4219 Further to IAUC 8082, R. Chassagne reports the discovery of an ap-

parent supernova (mag ∼ 15.1) on unfiltered images taken on Feb. 27.8 and Mar. 1.81 UT. SN 2003bj is located at α = 13^{h}18^{m}29^{s}.10, δ = −31^{o}37′37′′.7 (equinox 2000.0), which is 9′′.3 west and 13′′.7 north of the nucleus of IC 4219. Nothing was visible at this position on an image taken on 2002 Feb. 9.0 (limiting mag ∼ 18.0) or on a Digitized Sky Survey image taken on 1977 Mar. 20.7 (limiting mag 20.0).

SUPERNOVAE Further to their reports on IAUC 8084, A. V. Filippenko and R.

Chornock add that inspection of additional CCD spectra, also obtained on Feb. 27 or 28 UT, shows that SN 2003ac (IAUC 8064) is beginning to develop He I lines, as predicted on IAUC 8065. SN 2003am (IAUC 8069) is of type II in the nebular phase; Hα emission is very strong. SN 2003an (IAUC 8069) is of type Ia at redshift 0.037 (host galaxy), ∼ 2 weeks past maximum. SN 2003aq (IAUC 8072) is of type II, well into the plateau phase; Hα emission greatly dominates over the absorption component in the P-Cyg profile. SN 2003au (IAUC 8075) is of type Ia at redshift 0.031 (host galaxy), ∼ 1 month past maximum; the spectral features resemble those of the subluminous type-Ia SN 1991bg. SN 2003ay (IAUC 8079) is of type Ia at redshift 0.073 (from the supernova features), ∼ 2 weeks past maximum. The following redshifts were measured for objects reported on IAUC 8083: SN 2003V, 0.045 (host galaxy); SN 2003ah, 0.03 (supernova features); SN 2003av, 0.14 (supernova features); SN 2003ax, 0.054 (host galaxy); SN 2003bf, 0.034 (host galaxy).

2003aw P. A. Woudt and B. Warner, University of Cape Town, report that

high-speed CCD photometry of the object 2003aw at the South African Astronomical Observatory 1.0-m telescope yields a brightness (V ∼ 18.3 on Feb. 28.9 and Mar. 1.9 UT) that is modulated on a period of 2034 ± 3 s with a peak-to-peak amplitude of ∼ 0.1 mag, confirming the suggestion (IAUC 8084) that 2003aw is a double-degenerate helium-transferring binary currently in a high state. The observed periodicity is probably a superhump period, and shallow eclipses are evident.

2003 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green